OpenStudy (anonymous):

A runner takes 9 seconds to run a distance of 132 feet. What is the runner's speed in miles per hour?

OpenStudy (anonymous):

The speed = 132 ft/9 sec = 132*1/5280 miles/sec = 132*60/5280 miles/hr = 1.5 miles/hr